# Color it, Code it, Cancel it: $k$-local dynamical decoupling from classical additive codes

### v1

@@ III Permutation-invariant decoupling group from graph symmetry
Step 1: Define interaction hypergraph $I_{\rm Dev}$ Step 2: Color $I_{\rm Dev}$ with $\mathcal{C}[I_{\rm Dev}]$ Step 3: Construct symmetric hypergraph $I^{\prime}_{\rm Dev}$ Step 4: Form quotient graph $I^{\prime}_{\rm Dev}/\mathcal{C}[I_{\rm Dev}]$ Step 5: Generate universal sequence via additive code Step 6: Tailored sequence or Hamiltonian engineering
@@
Figure 2: Framework for constructing efficient dynamical decoupling sequences. (a) Flowchart of our approach. Starting from an interaction hypergraph, we color it, generate a symmetric variant, and construct its quotient. Additive codes are then used to produce efficient universal sequences, which are tailored to suppress specific Hamiltonian terms. (b) Example of each step in the protocol. We start from a simple 7-qubit loop (step 1) and we color it with a color partition under a nearest-neighbor interaction model (step 2). We then construct a color-preserving symmetric graph connecting all differnet colors (step 3), that permits to simplify the DD problem by just considering the corresponding quotient graph (step 4). The DD sequence is then constructed by using additive codes, allowing us to construct a universal decoupling group that freezes arbitrary two-local interactions (step 5). We then tailor the sequence for specific qubit architectures (step 6).
In this section, we systematically reformulate and extend the protocols introduced in Refs. [27, 24, 33] using the language of graph theory. This approach enables a more natural description of interactions in the Hamiltonian through hypergraphs and captures system symmetries via quotient graphs. The flowchart of the framework proposed in this work is summarized in Fig. 2.
@@ III.1 Hypergraph coloring
To describe a general Hamiltonian of $N$ interacting qubits, we introduce the interaction graph $I_{\rm Dev}=(V_{I},E_{I})$. The vertex set $V_{I}$ represents physical qubits. The interaction graph is a hypergraph, where each hyperedge connects a subset of vertices, potentially more than two. A hyperedge $e_{i}\in E_{I}$ connecting vertices $(v_{i_{1}},\dots,v_{i_{N}})$ corresponds to a general noise model in which arbitrary Pauli interactions from $\{I,X,Y,Z\}^{\otimes N}$ may jointly act on these qubits, unless otherwise specified. This corresponds to the first step of Fig. 2(a). A Pauli string associated with such a hyperedge is called $k$-local if it has support on exactly $k$ distinct vertices. Therefore, the interaction graph $I_{\rm Dev}$ characterizes a $k$-local Hamiltonian if and only if all of its hyperedges involve at most $k$ vertices. Under this definition, the Hamiltonian considered in Refs. [27, 24] are two-local and the interaction graph $I_{\rm Dev}$ coincides with the device connectivity graph.
@@
We now generalize the results of Refs. [27, 24] to arbitrary $k$-local Hamiltonians. In the first-order Magnus expansion given in Eq. (6), a vertex $v_{i}$ only interacts with other vertices that share a common hyperedge. This motivates a natural vertex coloring of the interaction graph. We assign a distinct color $c_{i}$ to each vertex $v_{i}\in V_{I}$, such that no two vertices $(v_{i},v_{j})\in E_{I}$ in a hyperedge share the same color (step 2 of Fig. 2(a)). We denote this graph coloring as $C[I_{\rm Dev}]$, and refer to the number of colors used (i.e. the chromatic number) as $\chi_{I}$.
@@
We remark that general hypergraph coloring problems can be addressed efficiently using well-established graph-coloring algorithms [40]. However, standard graph-coloring algorithms are typically designed for 2-local graphs, where each edge connects exactly two vertices. To apply these algorithms to a hypergraph coloring problem, we formally construct an auxiliary graph by adding an edge between each pair of vertices $v_{i}$ and $v_{j}$ that share a hyperedge in $I_{\rm Dev}$. A proper coloring of this auxiliary graph corresponds to a valid coloring of the original hypergraph.
@@
In the next section, we show that the color partition $C[I_{\rm Dev}]$ respects an underlying graph symmetry. Using this symmetry, the problem of designing a DD sequence for the interaction graph $I_{\rm Dev}$ is significantly simplified.
III.2 Symmetry and quotient graph
@@
We begin by constructing an expanded graph $I_{\rm Dev}^{{}^{\prime}}=(V_{I},E_{I}^{{}^{\prime}})$ from $I_{\rm Dev}$ such that $I_{\rm Dev}\subseteq I_{\rm Dev}^{{}^{\prime}}$, and the original coloring $C[I_{\rm Dev}]$ remains a valid vertex coloring for $I_{\rm Dev}^{{}^{\prime}}$. Because $I_{\rm Dev}\subseteq I_{\rm Dev}^{{}^{\prime}}$, any DD group $\mathcal{G}$ designed for $I_{\rm Dev}^{{}^{\prime}}$ is also valid decoupling group for $I_{\rm Dev}$
@@
More precisely, the expanded graph $I_{\rm Dev}^{{}^{\prime}}$ is constructed from $I_{\rm Dev}$ and its coloring $C[O_{\rm Dev}]$ as follows. The graph $I_{\rm Dev}^{{}^{\prime}}$ shares the same vertex set $V_{I}^{{}^{\prime}}=V_{I}$ as $I_{\rm Dev}$. We traverse the colored interaction graph $I_{\rm Dev}$ and record all tuples $(c_{i_{1}},\dots,c_{i_{N}})$ such that there exists a hyperedge in $I_{\rm Dev}$ connecting vertices with those colors. The new hyperedge set $E_{I^{\prime}}$ is then constructed by appending to the original set $E_{I}$ additional hyperedges of the form $(v_{i_{1}},\dots,v_{i_{N}})\notin E_{I}$, where the vertices $(v_{i_{1}},\dots,v_{i_{N}})$ have distinct colors that match one of the recorded color tuples $(c_{i_{1}},\dots,c_{i_{N}})$. This procedure ensures that the color partition $C[I_{\rm Dev}]$ remains valid for the expanded graph $I_{\rm Dev}^{{}^{\prime}}$, see step 3 of Fig. 2(a).
@@
To better illustrate this construction, here we apply it to an explicit example. We consider a system of seven qubits arranged in a bilinear array with two-local interactions. Figure 3(a) illustrates the corresponding interaction graph $I_{\rm Dev}$ along with its vertex coloring $C[I_{\rm Dev}]$. A general Hamiltonian for this system is
@@
where $\vec{h}_{i}$ are local fields and $A_{ij}$ are real coupling matrices that specify pairwise interactions. For this example, the expanded graph $I_{\rm Dev}^{{}^{\prime}}$ constructed from the above procedure is shown in Fig. 3(b).
@@
Figure 3: Colored bilinear array graph. (a) Interaction graph $I_{\rm Dev}$ of 77 physical qubits with two-local interactions arranged in a bilinear array, along with the minimal color partition $C[I_{\rm Dev}]$. (b) Expanded interaction graph $I_{\rm Dev}^{{}^{\prime}}$ constructed from $I_{\rm Dev}$ and $C[I_{\rm Dev}]$.
@@
The expanded graph $I^{{}^{\prime}}_{\rm Dev}$ may appear more complex than the original interaction graph $I_{\rm Dev}$, suggesting that designing a suitable decoupling group $\mathcal{G}$ for $I^{{}^{\prime}}_{\rm Dev}$ could be more challenging. However, $I^{{}^{\prime}}_{\rm Dev}$ inherits a strong symmetry from the color partition $C[I_{\rm Dev}]$. In particular, $I^{{}^{\prime}}_{\rm Dev}$ is invariant under permutations of vertices that share the same color, a symmetry we (with a slight abuse of notation) also denote by $C[I_{\rm Dev}]$. This invariance significantly simplifies the construction of the DD group: it implies that it suffices to consider DD groups $\mathcal{G}$ that themselves respect the symmetry $C[I_{\rm Dev}]$ [41]. Practically, this means applying identical control pulses to all qubits within the same color class.
@@
As a complementary view on this symmetry, rather than directly working with the symmetric decoupling group $\mathcal{G}$, we can instead analyze the quotient graph $I^{{}^{\prime}}_{\rm Dev}/C[I_{\rm Dev}]$. The quotient graph $I^{{}^{\prime}}_{\rm Dev}/C[I_{\rm Dev}]$ is formed by collapsing all vertices $v_{i}\in V_{I}^{{}^{\prime}}$ with the same color $c_{i}$ into a single super-vertex $c_{i}\in V_{I}^{{}^{\prime}}/C[I_{\rm Dev}]$ in the quotient graph. A hyperedge $(c_{i_{1}},\dots,c_{i_{N}})$ exists in the quotient graph if and only if there is at least one hyperedge $(v_{i_{1}},\dots,v_{i_{N}})\in E_{I}^{{}^{\prime}}$, such that the vertices $V_{i_{j}}$ are colored $c_{i_{j}}$ respectively. Figure 4 illustrates this transformation from the expanded graph $I_{\rm Dev}^{{}^{\prime}}$ (shown in Fig. 3) to its corresponding quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$. We emphasize that the same quotient graph also captures different topologies, including the ring topology in Fig. 2.
@@
It follows directly that if $\Pi_{\mathcal{G}}$ is a DD sequence that universally suppresses all interactions in the quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$, then this sequence can be lifted to a valid decoupling sequence for the expanded graph $I_{\rm Dev}^{{}^{\prime}}$, and consequently on the interaction graph $I_{\rm Dev}$. The lifting procedure consists of applying the same control pulses in parallel to all vertices in $I_{\rm Dev}^{{}^{\prime}}$ that share the same color $c_{i}$. This approach is valid for both bang-bang and bounded-strength control schemes.
@@
We note that the explicit construction of the expanded interaction graph $I_{\rm Dev}^{{}^{\prime}}$ in principle is not required to design the DD group. Rather, it provides a conceptual tool that reveals the underlying symmetry structure of the colored interaction graph. The expanded graph also provides a key general insight, namely any two interaction graphs that share the same expanded graph $I_{\rm Dev}^{{}^{\prime}}$ can be universally decoupled using the same DD group $\mathcal{G}$, see the examples in Figs. 2 and 3. In fact, an even stronger statement holds: if two interaction graphs share the same quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$ and interaction models, they admit the same DD group capable of universally suppressing their system interactions.
@@
Figure 4: Dynamical decoupling from orthogonal arrays. (a) The quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$ formed by collapsing all physical qubits of the same color in $I_{\rm Dev}^{{}^{\prime}}$ into a single equivalence class. (b) The orthogonal array $\text{OA}(L=16,\chi_{I}=3,S=4,k=2)$ which generates a universal DD sequence for two-local Hamiltonian defined by the quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$. Each column is colored according to the associated color qubits $c_{i}$.
@@
In the following section, we demonstrate how to obtain time-efficient decoupling groups for the quotient graph $I^{{}^{\prime}}_{\rm Dev}/C[G_{\rm Dev}]$ using constructions from classical coding theory.
@@
IV Time-efficient decoupling group from classical additive codes
We now discuss how to construct time-efficient DD groups using classical additive codes. Specifically, we establish a connection between DD sequences and classical additive codes by using orthogonal arrays [32, 30]. As a result, we can analyze the properties of a DD group in terms of its generators, in analogy to how stabilizer codes are characterized by their stabilizer generators [42].

### v2

@@ III Permutation-invariant decoupling group from hypergraph symmetry
Step 1: Define interaction hypergraph $I_{\rm Dev}$ Step 2: Color $I_{\rm Dev}$ with $\mathcal{C}[I_{\rm Dev}]$ Step 3: Construct symmetric hypergraph $I^{\prime}_{\rm Dev}$ Step 4: Form quotient hypergraph $I^{\prime}_{\rm Dev}/\mathcal{C}[I_{\rm Dev}]$ Step 5: Generate universal sequence via additive code Step 6: Tailored sequence or Hamiltonian engineering
@@
Figure 2: Framework for constructing efficient dynamical decoupling sequences. (a) Flowchart of our approach. Starting from an interaction hypergraph, we color it, generate a symmetric variant, and construct its quotient. Additive codes are then used to produce efficient universal sequences, which are tailored to suppress specific Hamiltonian terms. (b) Example of each step in the protocol. We start from a simple 7-qubit loop (step 1) and we color it assuming a nearest-neighbor interaction model (step 2). We then construct a color-preserving symmetric hypergraph connecting all different colors (step 3), which allows us to simplify the DD problem by just considering the corresponding quotient hypergraph (step 4). The DD sequence is then constructed by using additive codes, forming a universal decoupling group that suppresses arbitrary two-local interactions (step 5). We then tailor the sequence for specific qubit architectures (step 6).
In this section, we systematically reformulate and extend the ideas introduced in Refs. [brown2024efficient, evert2024syncopated, tong2024empirical] using the language of graph theory. This perspective enables a natural generalization of previous results from $2$-local interaction models to general $k$-local interaction models. In Sec. III.1, we describe the system Hamiltonian in terms of an interaction hypergraph and show that any $k$-local interaction induces a natural partition of physical qubits into equivalence color classes. In Sec. III, we demonstrate that this partition gives rise to a permutation symmetry that can be exploited to reduce the original interaction hypergraph to a simpler quotient hypergraph, by factoring out the symmetry group. As a result, the task of designing a DD sequence for the original interaction hypergraph is mapped to the simpler problem of designing a DD sequence for the quotient hypergraph.
@@ III.1 Hypergraph coloring
To describe a general Hamiltonian of $N$ interacting qubits, we introduce the interaction hypergraph $I_{\rm Dev}=(V_{I},E_{I})$. The vertex set $V_{I}=\{v_{1},\dots,v_{N}\}$ represents physical qubits. A hyperedge $e_{i}\in E_{I}$ connecting $k$ vertices $(v_{i_{1}},\dots,v_{i_{k}})$ corresponds to a general noise model in which arbitrary Pauli interactions from $\{I,X,Y,Z\}^{\otimes k}$ may jointly act on these $k$ qubits, unless otherwise specified. This corresponds to the first step of Fig. 2(a). A Pauli string associated with such a hyperedge is called $k$-local if it has support on exactly $k$ distinct vertices. Therefore, the interaction hypergraph $I_{\rm Dev}$ characterizes a $k$-local Hamiltonian if and only if all of its hyperedges involve at most $k$ vertices. The standard graph is simply 2-local hypergraph. Under this definition, the Hamiltonians considered in Refs. [brown2024efficient, evert2024syncopated] are 2-local and the interaction hypergraph $I_{\rm Dev}$ coincides with the device connectivity graph.
@@
We now generalize the results of Refs. [brown2024efficient, evert2024syncopated] to arbitrary $k$-local Hamiltonians. In the first-order Magnus expansion given in Eq. (6), a vertex $v_{i}$ only interacts with other vertices that share a common hyperedge. This motivates a natural partition of the interaction hypergraph $I_{\rm Dev}$. We assign a distinct color $c_{i}$ to each vertex $v_{i}\in V_{I}$, such that no two vertices $v_{i},v_{j}$ in a hyperedge share the same color (step 2 of Fig. 2(a)). We denote this hypergraph coloring as $C[I_{\rm Dev}]$, and refer to the number of colors used (i.e. the chromatic number) as $\chi_{I}$.
@@
We remark that this hypergraph coloring problem can be addressed efficiently using well-established graph-coloring algorithms [mostafaie2020systematic]. However, standard graph-coloring algorithms are typically designed for graphs and not hypergraphs. To apply these algorithms to a hypergraph coloring problem, we formally construct an auxiliary graph by adding an edge between each pair of vertices $v_{i}$ and $v_{j}$ that share a hyperedge in $I_{\rm Dev}$. A proper coloring of this auxiliary graph corresponds to a valid coloring of the original hypergraph $I_{\rm Dev}$..
@@
In the next section, we show that the color partition $C[I_{\rm Dev}]$ respects an underlying hypergraph symmetry. Using this symmetry, the problem of designing a DD sequence for the interaction hypergraph $I_{\rm Dev}$ is significantly simplified.
III.2 Symmetry and quotient hypergraph
@@
We now show how the color partition $C[I_{\rm Dev}]$ can be exploited to simplify the problem. The strategy is as follows. We first construct an expanded hypergraph $I_{\rm Dev}^{{}^{\prime}}=(V_{I},E_{I}^{{}^{\prime}})$ from $I_{\rm Dev}$ such that $I_{\rm Dev}\subseteq I_{\rm Dev}^{{}^{\prime}}$, and the color partition $C[I_{\rm Dev}]$ remains a valid vertex coloring for $I_{\rm Dev}^{{}^{\prime}}$. Because $I_{\rm Dev}\subseteq I_{\rm Dev}^{{}^{\prime}}$, any decoupling group $\mathcal{G}$ designed for $I_{\rm Dev}^{{}^{\prime}}$ is also valid decoupling group for $I_{\rm Dev}$. We then show that designing a decoupling group for $I_{\rm Dev}^{{}^{\prime}}$ is equivalent to designing one for the significantly simpler quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$, which contains only $\chi_{I}$ vertices. In this way, the original problem is reduced to constructing a decoupling group for the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$.
@@
More precisely, the expanded hypergraph $I_{\rm Dev}^{{}^{\prime}}$ is constructed from $I_{\rm Dev}$ and its coloring $C[I_{\rm Dev}]$ as follows. The hypergraph $I_{\rm Dev}^{{}^{\prime}}$ shares the same vertex set $V_{I}^{{}^{\prime}}=V_{I}$ as $I_{\rm Dev}$. We traverse the colored hypergraph $I_{\rm Dev}$ and record all tuples $(c_{i_{1}},\dots,c_{i_{k}})$ such that there exists a hyperedge in $I_{\rm Dev}$ connecting vertices with those colors. The new hyperedge set $E_{I}^{\prime}$ is then constructed by appending to the original set $E_{I}$ additional hyperedges of the form $(v_{i_{1}},\dots,v_{i_{k}})\notin E_{I}$, where the vertices $(v_{i_{1}},\dots,v_{i_{k}})$ have distinct colors that match one of the recorded color tuples $(c_{i_{1}},\dots,c_{i_{k}})$. This procedure ensures that the color partition $C[I_{\rm Dev}]$ remains valid for the hypergraph $I_{\rm Dev}^{{}^{\prime}}$, see step 3 of Fig. 2(a).
@@
To better illustrate this construction, here we apply it to an explicit example. We consider a system of seven qubits arranged in a bilinear array with two-local interactions. Figure 3(a) illustrates the corresponding interaction hypergraph $I_{\rm Dev}$ along with its vertex coloring $C[I_{\rm Dev}]$. A general Hamiltonian for this system is
@@
where $\vec{h}_{i}$ are local fields and $A_{ij}$ are real coupling matrices that specify pairwise interactions. For this example, the expanded hypergraph $I_{\rm Dev}^{{}^{\prime}}$ constructed from the above procedure is shown in Fig. 3(b), with new edges connecting vertices of different colors.
@@
Figure 3: Colored bilinear array graph. (a) Interaction hypergraph $I_{\rm Dev}$ of $7$ physical qubits with two-local interactions arranged in a bilinear array, along with the minimal color partition $C[I_{\rm Dev}]$. (b) Expanded hypergraph $I_{\rm Dev}^{{}^{\prime}}$ constructed from $I_{\rm Dev}$ and $C[I_{\rm Dev}]$.
@@
The expanded hypergraph $I^{{}^{\prime}}_{\rm Dev}$ may appear more complex than the original interaction hypergraph $I_{\rm Dev}$, suggesting that designing a suitable decoupling group $\mathcal{G}$ for $I^{{}^{\prime}}_{\rm Dev}$ could be more challenging. However, $I^{{}^{\prime}}_{\rm Dev}$ inherits a strong symmetry from the color partition $C[I_{\rm Dev}]$. In particular, $I^{{}^{\prime}}_{\rm Dev}$ is invariant under permutations of vertices that share the same color, a symmetry we (with a slight abuse of notation) also denote by $C[I_{\rm Dev}]$. This invariance significantly simplifies the construction of the decoupling group: it implies that it suffices to consider decoupling groups $\mathcal{G}$ that themselves respect the symmetry $C[I_{\rm Dev}]$ [zanardi1999symmetrizing]. Practically, this means applying identical control pulses to all qubits within the same color class.
@@
As a complementary view on this symmetry, rather than directly working with the symmetric decoupling group $\mathcal{G}$, we can instead analyze the quotient hypergraph $I^{{}^{\prime}}_{\rm Dev}/C[I_{\rm Dev}]$. The quotient hypergraph $I^{{}^{\prime}}_{\rm Dev}/C[I_{\rm Dev}]$ is formed by collapsing all vertices $v_{i}\in V_{I}^{{}^{\prime}}$ with the same color $c_{i}$ into a single super-vertex $c_{i}\in V_{I}^{{}^{\prime}}/C[I_{\rm Dev}]$ in the quotient hypergraph. We will refer to the super-vertex $c_{i}$ in the quotient hypergraph $I^{{}^{\prime}}_{\rm Dev}/C[I_{\rm Dev}]$ as the color class $c_{i}$. A hyperedge $(c_{i_{1}},\dots,c_{i_{k}})$ exists in $I^{{}^{\prime}}_{\rm Dev}/C[I_{\rm Dev}]$ if and only if there is at least one hyperedge $(v_{i_{1}},\dots,v_{i_{k}})\in E_{I}^{{}^{\prime}}$, such that the vertices $v_{i_{j}}$ are colored $c_{i_{j}}$ respectively. Figure 4 illustrates this transformation from the expanded hypergraph $I_{\rm Dev}^{{}^{\prime}}$ (shown in Fig. 3) to its corresponding quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$. We emphasize that the same quotient hypergraph also captures different topologies, including the ring topology in Fig. 2.
@@
It follows directly that if there is a DD sequence that universally suppresses all interactions in the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$, then this sequence can be lifted to a valid decoupling sequence for the expanded hypergraph $I_{\rm Dev}^{{}^{\prime}}$, and consequently on the interaction hypergraph $I_{\rm Dev}$. The lifting procedure consists of applying the same control pulses in parallel to all vertices in $I_{\rm Dev}^{{}^{\prime}}$ that share the same color $c_{i}$. This approach is valid for both bang-bang and bounded-strength control schemes.
@@
We note that the explicit construction of the expanded interaction hypergraph $I_{\rm Dev}^{{}^{\prime}}$ in principle is not required to design the DD group. Rather, it provides a conceptual tool that reveals the underlying symmetry structure of the colored interaction hypergraph. The expanded hypergraph also provides a key general insight, namely any two interaction hypergraphs that share the same expanded hypergraph $I_{\rm Dev}^{{}^{\prime}}$ can be universally decoupled using the same decoupling group $\mathcal{G}$, see the examples in Figs. 2 and 3. In fact, an even stronger statement holds: if two interaction hypergraphs share the same quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$ and interaction model, they admit the same decoupling group capable of universally suppressing their system interactions.
@@
Figure 4: Dynamical decoupling from orthogonal arrays. (a) The quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$ formed by collapsing all physical qubits of the same color in $I_{\rm Dev}^{{}^{\prime}}$ into a single equivalence class. (b) The orthogonal array $\text{OA}(L=16,\chi_{I}=3,S=4,k=2)$ which generates a universal DD sequence for two-local Hamiltonian defined by the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$. Each column is colored according to the associated color class $c_{i}$.
@@
In the following section, we demonstrate how to obtain time-efficient decoupling groups for the quotient hypergraph $I^{{}^{\prime}}_{\rm Dev}/C[G_{\rm Dev}]$ using constructions from classical coding theory.
@@
IV Time-efficient decoupling groups from classical additive codes
In the previous section, we demonstrated how the problem of constructing a decoupling group $\mathcal{G}$ for an arbitrary $k$-local interaction hypergraph $I_{\rm Dev}$ can be reduced to the simpler task of constructing a decoupling group for the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$. In this section, we explain how to build efficient decoupling groups for $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$ using tools from classical coding theory. In Sec. IV.1, we introduce the connection between decoupling groups $\mathcal{G}$ and a combinatorial object known as orthogonal arrays as found in previous works [Rotteler_2006, Bookatz_2016]. These objects enable constructions of decoupling groups that achieve favorable trade-offs between the DD sequence length $L$ and the chromatic number $\chi_{I}$. In Sec. IV.2, we show how to construct the required orthogonal arrays from classical additive codes. In Sec. IV.3, we further exploit the structure of classical additive codes to show that the decoupling group can be completely characterized by its generators, thereby reducing the design of the full decoupling group to the simpler task of designing its generators. Finally, in Sec. IV.4, we present a simple numerical algorithm for finding decoupling groups that selectively cancel certain interactions while preserving others.